OpenStudy (anonymous):

What is the equation of the graphed line? Hint: Determine the slope of the line. y = x

OpenStudy (anonymous):

Heres the graph:

OpenStudy (igreen):

Let's take any two points on the line. Let's take (-1, 5) and (0, 0). We can plug these into the slope formula. \(m =\dfrac{y_2-y_1}{x_2-x_1}\) \(m =\dfrac{0-5}{0+1}\) \(m =\dfrac{-5}{1}\) Can you divide that? @Leviatho

OpenStudy (anonymous):

-5 right.

OpenStudy (anonymous):

Find the (x,y) coordinates to do the questions. Simple coordinates: like (-1,5) and (1,-5) for slope m = (y1-y)/(x1-x) m = (-5-5)/(1-(-1))) = -10/2 now to find the equation of line use point slope formula \[(y-y_1) = m(x-x_1)\] put any one of the coordinates in place of (x1,y1)

OpenStudy (igreen):

Yep, so the slope is -5.

OpenStudy (igreen):

Do you know what the y-intercept is? @Leviatho

OpenStudy (anonymous):

I dont think so..

OpenStudy (igreen):

The y-intercept is the point where the line crosses on the y-axis.

OpenStudy (igreen):

Can you see where that is on the graph? @Leviatho

OpenStudy (anonymous):

Would that be 0?

OpenStudy (igreen):

Yep, you got it! And we plug these into the equation. y = mx + b Where m = slope, and b = y-intercept. We found out that slope = -5, and y-intercept = 0 y = -5x + 0 Simplify: y = -5x So that's your equation!

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Leviatho

OpenStudy (anonymous):

Thanks!

OpenStudy (igreen):

No problem!